# FIARSE: Model-Heterogeneous Federated Learning via Importance-Aware Submodel Extraction

### v1

@@ Abstract
In federated learning (FL), accommodating clients’ varied computational capacities poses a challenge, often limiting the participation of those with constrained resources in global model training. To address this issue, the concept of model heterogeneity through submodel extraction has emerged, offering a tailored solution that aligns the model’s complexity with each client’s computational capacity. In this work, we propose Federated Importance-Aware Submodel Extraction (FIARSE), a novel approach that dynamically adjusts submodels based on the importance of model parameters, thereby overcoming the limitations of previous static and dynamic submodel extraction methods. Compared to existing works, the proposed method offers a theoretical foundation for the submodel extraction and eliminates the need for additional information beyond the model parameters themselves to determine parameter importance, significantly reducing the overhead on clients. Extensive experiments are conducted on various datasets to showcase superior performance of the proposed FIARSE.
@@ 1 Introduction
Federated learning (FL) [29, 43] stands out as a promising distributed training paradigm, in which the clients enjoy mutual information without jeopardizing data privacy. Specifically, the FL server requests the clients to train a model with their local data and aggregates the models into a global one. Such a paradigm, however, may fail in a real-world FL system, where the clients usually have varying computation capacities [35, 11, 38], likely preventing the clients with insufficient computation resources from being involved in training a large global model.
@@
Depending on whether the submodels undergo reconstruction during the model training process, existing works fall into two categories: static submodel extraction [11, 16, 28, 26, 54, 18, 59, 47] and dynamic submodel extraction [2, 37, 7].
@@ 3.2 Algorithm Description
Partial client participation is one of the features of FL algorithms because the server is unlikely to handle all the communications from all clients, especially when the number of clients is considerably large [36, 57, 25]. Therefore, we present our algorithm FIARSE in support of partial client participation: At the beginning of each communication round $t\in\{0,1,\dots\}$, the server uniformly samples a group of clients from $[N]$ without replacement, denoted by $\mathcal{A}$, which consists of $A$ clients. Subsequently, the server broadcasts the submodels to the selected clients and collects and merges their updates into the global model. In the rest of the section, we comprehensively discuss the details of these steps and exemplify them with $t$-th round.
@@ 4 Convergence Analysis
Lipschitz-smooth assumption has gained widespread acceptance in machine learning research, as evidenced by its incorporation in various studies such as [1, 63, 42, 6, 57, 58]. Assumption 4.1 extends this assumption to a scenario where a binary mask is calculated based on the model parameters and subsequently applied to the model. The second assumption, widely made in the previous FL studies [57, 58], establishes bounds between the local objectives and the global objective due to the occurrence of non-i.i.d. data. Assumption 4.3 draws inspiration from [63, 42] and characterizes the masking performance by comparing gradients with and without the application of the mask. Notably, if $\mathcal{M}^{(i)}_{t}=\boldsymbol{1}^{d}$ (i.e., setting the threshold for $\theta_{i}=0$), then $\delta_{t}^{2}=0$.
@@
Suppose that Assumption 4.1, 4.2 and 4.3 hold. We define $F(\tilde{\boldsymbol{x}})\overset{\triangle}{=}F\left(\tilde{\boldsymbol{x}},\boldsymbol{1}^{N\times d}\right)$, and $F(\tilde{\boldsymbol{x}})\geq F_{*}$ for all $\tilde{\boldsymbol{x}}\in\mathbb{R}^{d}$. Let the local learning rate satisfy
@@
Suppose that Assumption 4.1, 4.2 and 4.3 hold. We define $F(\tilde{\boldsymbol{x}})\overset{\triangle}{=}F\left(\tilde{\boldsymbol{x}},\boldsymbol{1}^{N\times d}\right)$, and $F(\tilde{\boldsymbol{x}})\geq F_{*}$ for all $\tilde{\boldsymbol{x}}\in\mathbb{R}^{d}$. Let the local learning rate $\eta_{l}=\frac{1}{K\sqrt{T}}$, and the global learning rate $\eta_{s}=1$, where $T$ is the total communication rounds. Then, under non-convex objectives, FIARSE converges to a small neighborhood of a stationary point of standard FL as $T$ is large enough, i.e.,
@@
Regarding the first term on the right-hand side (RHS) of Equation (5), it can approach zero as $T$ tends to be infinite. However, an intriguing question arises concerning the potential for the second term to reach zero, given that the norm $\left\|\tilde{\boldsymbol{x}}_{t}\right\|_{2}^{2}$ cannot be zero. In Assumption 4.3, an obvious case that $\delta_{t}$ is always zero is when all clients use the full model, i.e., the threshold $\theta_{i}$ in Equation (3) is set to 0 in our case. Therefore, the theorem is consistent with the state-of-the-art works [27, 57] in terms of the convergence rate, i.e., $O\left(1/\sqrt{T}\right)$, under model-homogeneous FL. As for model-heterogeneity FL, it is noteworthy to consider that $\delta_{t}$ may tend towards zero as $t\rightarrow\infty$, leading to the outcome where $\frac{1}{T}\sum_{t\in[T]}\delta_{t}^{2}\left\|\tilde{\boldsymbol{x}}_{t}\right\|_{2}^{2}$ approaches zero. As explored by [63, 42, 55], this conclusion holds because the submodel could replicate the outcomes of the full model on all clients after substantial communication rounds $T$. To sum up, our proposed algorithm can achieve a convergence rate of $O\left(\frac{N\sigma^{2}}{A\sqrt{T}}\right)$, as long as $\lim_{t\rightarrow\infty}\delta_{t}=0$.
Regardless of the constant terms, the proposed FIARSE converges at a rate of $O\left(\sigma^{2}/\sqrt{T}\right)$ under full-client participation (i.e., $A=N$). Recently, some works have provided their convergence rates under model-heterogeneous FL with full-client participation, e.g., pruning-greedy [63] with $O(M\sigma^{2}/\sqrt{T})$ and FedDSE [50] with $O(K\sigma^{2}/\sqrt{T})$. Apparently, our proposed method can perform better from the theoretical perspective.
@@ 5 Experiments
Table 1: Test accuracy under four different submodel sizes. To be more specific, the columns from “Local” to “Model (1.0)” evaluate the test accuracy on the local test datasets, while “Global” evaluates the test accuracy of the global model on the global test dataset.
@@
(a) histogram of HeteroFL
(b) histogram of FedRolex
(c) histogram of ScaleFL
(d) histogram of FIARSE
Figure 2: Histograms of various submodel extraction methods on CIFAR-10 dataset. Each histogram illustrates the numbers of clients along with their respective performance in terms of Accuracy.
5.1 Setup
Datasets and Models.
Data Heterogeneity.
For CIFAR-10 and CIFAR-100, we follow [17, 23] and partition the datasets into 100 clients based on a Dirichlet distribution setting $\alpha=0.3$. As for AGNews, we partition the datasets for 200 clients with Dirichlet distribution as well, but it is with the parameter of $\alpha=1.0$. Note that no public dataset is used by server or clients during the training stage. In the testing phase, we refer to the superset of all clients’ test datasets as “global test dataset".
System Heterogeneity.
Implementation.
In this setting, we set the participation ratio to 10% by default. We perform 800-round training on the CV tasks while running for 400 rounds on the NLP task. To avoid the randomness of the results, we averaged the results from three different random seeds. In the experiments, we report the results of all the baselines based on the best hyperparameter settings. Due to the space limitation, more experimental results and analysis are deferred to Appendix D.
@@ 5.2 Submodel Performance on Local Dataset
In this setting, we evaluate the submodels’ performance on each client’s test datasets. To be specific, the local models are extracted from the global models. Figure 2 comprehensively illustrates the number of clients across different test accuracies. Among the four figures presented, all exhibit a left-skewed distribution, with the exception of FedRolex. This outcome aligns with our expectations, as FedRolex employs a rolling-based approach and is unable to concentrate on optimizing submodel performance on the local dataset, while the rest three approaches can spare efforts on a specific (HeteroFL and ScaleFL) or important (FIARSE) part, effectively addressing performance on local datasets. Among these three approaches, we notice that FIARSE stands out with the best results, showcasing superior performance as more clients achieve higher accuracy compared to the other two methods. The averaged results are also reported by Table 1 under the column of CIFAR-10 and “Local” to “Model (1.0)”. Specifically, “Model (1/64)” to “Model (1.0)” shows the averaged local performances classified by the model sizes, and the “Local” shows the result by averaging across these four columns. Table 1 also presents the test accuracy of CIFAR-100 and AGNews. The proposed FIARSE achieves at least 2% better than other baselines under these datasets.
@@ 5.3 Submodel Performance on Global Dataset
(a) Model size 1/64
(b) Model size 1/16
(c) Model size 1/4
(d) Model size 1.0
Figure 3: Performance changes of various strategies for submodel extraction in terms of the test accuracy versus the communication round on global test dataset of CIFAR-10.
In this setting, we evaluate the performance of submodels with various sizes on the global test dataset to assess the generalizability of our proposed algorithm. Table 1 presents the results of two CV datasets and one NLP dataset under the column “Global”. In conjunction with Figure 3, our proposed method FIARSE constantly outperforms other baselines in all submodels with different sizes by a substantial margin. Figure 3 dives into the details of training and shows the test accuracy trend throughout the communication rounds. Consider the submodels are expected to surpass a 70% test accuracy threshold. As previously discussed, FIARSE ultimately achieves superior test accuracy compared to other baselines. Across model sizes of {1/64, 1/16, 1/4}, our proposed method requires fewer rounds to reach the targeted accuracy compared to other baselines. While the performance disparity between FIARSE and FedRolex is less discernible under the full model (Model (1.0)), both methods significantly outpace static submodel extraction approaches in achieving 70% accuracy. In summary, the proposed method stands out by attaining the desired submodels with the fewest rounds among the approaches implemented in this section.
Figure 4: Performance changes of different submodel extraction methods to various submodel sizes.
@@ 5.4 Unparticipated Clients Performance
The above evaluations are conducted on the clients who participated in the training process. However, a more general scenario includes clients who skip the training phase but need models to process newly arrived data. In such cases, our algorithm can enable the server to customize models for them from the trained global model. Same as the expression in Line 3 of Algorithm 1, the server extracts a submodel encompassing the parameters $|\tilde{\boldsymbol{x}}_{t}|\geq\theta_{i}$ and sends it to the client. Note that the unparticipated clients could have capacities different from that of any client involved in the training. In Figure 4, we employ our algorithm as well as the baselines to extract submodels with different sizes and compare their performance on unparticipated clients. Generally, the performances of all methods increase as the size of extracted submodels grows. Our algorithm consistently outperforms the baseline methods. In contrast, alternative approaches, particularly static ones like HeteroFL and ScaleFL, suffer from more significant drops in performance. For the case of extracting a submodel that is much smaller than the minimum size involved in the training stage, our proposed method demonstrates remarkable superiority, outperforming existing works by at least 10%.
@@ 6 Conclusion
This work introduces an algorithm for model-heterogeneity FL, named FIARSE, that utilizes importance-aware operation to extract various sizes of submodels. In detail, we utilize TCB-GD that is able to optimize the clients’ local parameters to reflect their importance levels. Subsequently, we provide a theoretical analysis and highlight that the proposed work can converge to a stationary point at the rate of $O\left(1/\sqrt{T}\right)$, where $T$ is the number of communication rounds. Extensive experiments are conducted on ResNet-18 and Roberta-base that demonstrate the significant superiority of our proposed method over the state-of-the-art works.
@@
This work addresses the challenge of model-heterogeneous federated learning, where the computation capacities vary among the clients. It is suitable for real-world deployment and improves the efficiency of on-device training.

### v2

@@ Abstract
In federated learning (FL), accommodating clients’ varied computational capacities poses a challenge, often limiting the participation of those with constrained resources in global model training. To address this issue, the concept of model heterogeneity through submodel extraction has emerged, offering a tailored solution that aligns the model’s complexity with each client’s computational capacity. In this work, we propose Federated Importance-Aware Submodel Extraction (FIARSE), a novel approach that dynamically adjusts submodels based on the importance of model parameters, thereby overcoming the limitations of previous static and dynamic submodel extraction methods. Compared to existing works, the proposed method offers a theoretical foundation for the submodel extraction and eliminates the need for additional information beyond the model parameters themselves to determine parameter importance, significantly reducing the overhead on clients. Extensive experiments are conducted on various datasets to showcase the superior performance of the proposed FIARSE.
@@ 1 Introduction
Federated learning (FL) [36, 52] stands out as a promising distributed training paradigm, in which the clients enjoy mutual information without jeopardizing data privacy. Specifically, the FL server requests the clients to train a model with their local data and aggregates the models into a global one. Such a paradigm, however, may fail in a real-world FL system, where the clients usually have varying computation capacities [42, 13, 46], likely preventing the clients with insufficient computation resources from being involved in training a large global model [73].
@@
Depending on whether the submodels undergo reconstruction during the model training process, existing works fall into two categories: static submodel extraction [13, 35, 33, 67, 25, 76, 58] and dynamic submodel extraction [7, 21, 2, 45, 8].
@@
2 Related Work
This section discusses the state-of-the-art works that are most relevant to our research. Appendix A provides a more comprehensive review.
Computation Heterogeneity in FL.
Computation heterogeneity in FL refers to the varying computational capacities among clients, including differences in hardware capabilities and resource availability. One typical solution is to allow faster clients to perform more local updates, while slower ones update their local models fewer times [64, 43, 53, 50, 57, 72]. However, these approaches require clients to train the full model, which becomes infeasible when some clients cannot load the full model due to limited computation resources. In our work, we extract a submodel for each client that fits within their computational capacity.
Model Customization in FL.
Model customization allows the clients to build their local models that align with their local computation resources [46, 77, 28, 11, 66, 80]. To aggregate these heterogeneous models, the method often employs distillation for knowledge transfer, which requires a shared public dataset. However, this approach becomes infeasible when a shared public dataset is unavailable [59, 3]. In our work, we eliminate the need for a public dataset, broadening its applicability to a wider range of scenarios.
Model Sparsification in FL.
Model sparsification, also known as model pruning, removes the unimportant model parameters from a deep learning model, reducing computation overhead and tailoring model sizes to suit clients with varying computational resources [81, 45, 8]. For example, Flado [45] achieves model sparsification by requiring each client to maintain the importance levels of model parameters and extracting a submodel that encompasses the most important model parameters. While effective, this method incurs considerable overhead on each client in terms of both storage and computation, posing significant challenges for resource-constrained devices. In contrast, our work implicitly represents parameter importance through their values, eliminating the need to maintain separate importance scores for each parameter.
@@ 4.2 Algorithm Description
Partial client participation is one of the features of FL algorithms because the server is unlikely to handle all the communications from all clients, especially when the number of clients is considerably large [44, 74, 32, 71, 10]. Therefore, we present our algorithm FIARSE in support of partial client participation: At the beginning of each communication round $t\in\{0,1,\dots\}$, the server uniformly samples a group of clients from $[N]$ without replacement, denoted by $\mathcal{A}$, which consists of $A$ clients. Subsequently, the server broadcasts the submodels to the selected clients and collects and merges their updates into the global model. In the rest of the section, we comprehensively discuss the details of these steps and exemplify them with $t$-th round.
@@ 5 Convergence Analysis
Lipschitz-smooth assumption has gained widespread acceptance in machine learning research, as evidenced by its incorporation in various studies such as [1, 81, 51, 6, 74, 75, 70]. Assumption 5.1 extends this assumption to a scenario where a binary mask is calculated based on the model parameters and subsequently applied to the model. The second assumption, widely made in the previous FL studies [74, 75], establishes bounds between the local objectives and the global objective due to the occurrence of non-i.i.d. data. Assumption 5.3 draws inspiration from [81, 51] and characterizes the masking performance by comparing gradients with and without the application of the mask. Notably, if $\mathcal{M}^{(i)}_{t}=\boldsymbol{1}^{d}$ (i.e., setting the threshold for $\theta_{i}=0$), then $\delta_{t}^{2}=0$.
@@
Suppose that Assumption 5.1, 5.2 and 5.3 hold. We define $F(\tilde{\boldsymbol{x}})\overset{\triangle}{=}F\left(\tilde{\boldsymbol{x}},\boldsymbol{1}^{N\times d}\right)$, and $F(\tilde{\boldsymbol{x}})\geq F_{*}$ for all $\tilde{\boldsymbol{x}}\in\mathbb{R}^{d}$. Let the local learning rate satisfy
@@
Suppose that Assumption 5.1, 5.2 and 5.3 hold. We define $F(\tilde{\boldsymbol{x}})\overset{\triangle}{=}F\left(\tilde{\boldsymbol{x}},\boldsymbol{1}^{N\times d}\right)$, and $F(\tilde{\boldsymbol{x}})\geq F_{*}$ for all $\tilde{\boldsymbol{x}}\in\mathbb{R}^{d}$. Let the local learning rate $\eta_{l}=\frac{1}{K\sqrt{T}}$, and the global learning rate $\eta_{s}=1$, where $T$ is the total communication rounds. Then, under non-convex objectives, FIARSE converges to a small neighborhood of a stationary point of standard FL as $T$ is large enough, i.e.,
@@
Regarding the first term on the right-hand side (RHS) of Equation (5), it approaches zero as $T$ tends to infinity. However, an intriguing question arises concerning the potential for the second term to reach zero, given that the norm $\left\|\tilde{\boldsymbol{x}}_{t}\right\|_{2}^{2}$ cannot be zero. According to Assumption 5.3, a straightforward case where $\delta_{t}$ is always zero occurs when all clients use the full model – that is, when the threshold $\theta_{i}$ in equation Equation (3) is set to zero in the proposed FIARSE. Therefore, under model-homogeneous federated learning (FL), our theorem aligns with state-of-the-art works [34, 74] in terms of the convergence rate, which focuses solely on the number of communication rounds $T$, i.e., $O\left(1/\sqrt{T}\right)$.
For model-heterogeneous FL, it is noteworthy that $\delta_{t}$ may tend toward zero as $t\rightarrow\infty$, leading to $\frac{1}{T}\sum_{t=1}^{T}\delta_{t}^{2}\left\|\tilde{\boldsymbol{x}}_{t}\right\|_{2}^{2}$ approaching zero. As explored by [81, 51, 69], this occurs because the submodels can replicate the performance of the full model on all clients after a substantial number of communication rounds $T$. Consequently, our proposed algorithm can achieve a convergence rate of $O\left(\frac{N\sigma^{2}}{A\sqrt{T}}\right)$, as long as $\lim_{t\rightarrow\infty}\delta_{t}=0$.
Ignoring constant terms, the proposed FIARSE converges at a rate of $O\left(\sigma^{2}/\sqrt{T}\right)$ under full-client participation (i.e., $A=N$). Recently, some works have reported their convergence rates under model-heterogeneous FL with full-client participation, such as pruning-greedy [81] with a rate of $O\left(M\sigma^{2}/\sqrt{T}\right)$ and FedDSE [62] with $O\left(K\sigma^{2}/\sqrt{T}\right)$. Evidently, our proposed method exhibits better theoretical performance.
@@ 6 Experiments
6.1 Setup
Datasets and Models.
Data Heterogeneity.
For CIFAR-10 and CIFAR-100, we follow [22, 30] and partition the datasets into 100 clients based on a Dirichlet distribution setting $\alpha=0.3$. As for AGNews, we partition the datasets for 200 clients with Dirichlet distribution as well, but it is with the parameter of $\alpha=1.0$. Note that the server or clients do not use any public datasets during the training stage. In the testing phase, we refer to the superset of all clients’ test datasets as "global test dataset".
System Heterogeneity.
Implementation.
In this setting, we set the participation ratio to 10% by default. We perform 800-round training on the CV tasks while running for 300 rounds on the NLP task. To avoid the randomness of the results, we averaged the results from three different random seeds. In the experiments, we report the results of all the baselines based on the best hyperparameter settings. Due to the space limitation, more experimental results and analysis are deferred to Appendix D. Our code is released at https://github.com/HarliWu/FIARSE.
Table 1: Test accuracy under four different submodel sizes. To be more specific, the columns from "Local" to "Model (1.0)" evaluate the test accuracy on the local test datasets, while "Global" evaluates the test accuracy of the global model on the global test dataset.
@@ 6.1 Setup
(a) HeteroFL Histogram (b) FedRolex Histogram (c) ScaleFL Histogram (d) FIARSE Histogram
Figure 2: Histograms of various submodel extraction methods on CIFAR-10 under four submodel sizes. Each histogram shows the number of clients achieving different levels of test accuracy.
@@ 6.2 Submodel Performance on Local Dataset
In this setting, we evaluate the submodels’ performance on each client’s test datasets. To be specific, the local models are extracted from the global models. Figure 2 comprehensively illustrates the number of clients across different test accuracies. Among the four figures presented, all exhibit a left-skewed distribution, with the exception of FedRolex. This outcome aligns with our expectations, as FedRolex employs a rolling-based approach and is unable to concentrate on optimizing submodel performance on the local dataset, while the rest three approaches can spare efforts on a specific (HeteroFL and ScaleFL) or important (FIARSE) part, effectively addressing performance on local datasets. Among these three approaches, we notice that FIARSE stands out with the best results, showcasing superior performance as more clients achieve higher accuracy compared to the other two methods. The averaged results are also reported by Table 1 under the column of CIFAR-10 and "Local" to "Model (1.0)". Specifically, "Model (1/64)" to "Model (1.0)" shows the averaged local performances classified by the model sizes, and the "Local" shows the result by averaging across these four columns. Table 1 also presents the test accuracy of CIFAR-100 and AGNews. The proposed FIARSE achieves at least 2% better than other baselines under these datasets.
@@ 6.3 Submodel Performance on Global Dataset
(a) Model Size 1/64 (b) Model Size 1/16 (c) Model Size 1/4 (d) Model Size 1.0
(e) Model Size 1/64 (f) Model Size 1/16 (g) Model Size 1/4 (h) Model Size 1.0
Figure 3: Comparison of test accuracy across communication rounds for different submodel extraction strategies under four varying model sizes (1/64, 1/16, 1/4, 1/0) on global test datasets of CIFAR-10 (upper, a – d) and CIFAR-100 (lower, e – h).
In this setting, we evaluate the performance of submodels with various sizes on the global test dataset to assess the generalizability of our proposed algorithm. Table 1 presents the results of two CV datasets and one NLP dataset under the column "Global". In conjunction with Figure 3, our proposed method FIARSE constantly outperforms other baselines in all submodels with different sizes by a substantial margin. Figure 3 dives into the details of training and shows the test accuracy trend throughout the communication rounds. Consider the submodels are expected to surpass a 70% test accuracy threshold. As previously discussed, FIARSE ultimately achieves superior test accuracy compared to other baselines. Across model sizes of {1/64, 1/16, 1/4}, our proposed method requires fewer rounds to reach the targeted accuracy compared to other baselines. While the performance disparity between FIARSE and FedRolex is less discernible under the full model (Model (1.0)), both methods significantly outpace static submodel extraction approaches in achieving 70% accuracy. In summary, the proposed method stands out by attaining the desired submodels with the fewest rounds among the approaches implemented in this section.
@@ 6.4 Unparticipated Clients Performance
The above evaluations are conducted on the clients who participated in the training process. However, a more general scenario includes clients who skip the training phase but need models to process newly arrived data. In such cases, our algorithm can enable the server to customize models from the trained global model for them. Same as the expression in Line 3 of Algorithm 1, the server extracts a submodel encompassing the parameters $|\tilde{\boldsymbol{x}}_{t}|\geq\theta_{i}$ and sends it to the client. Note that the unparticipated clients could have capacities different from that of any client involved in the training.
Figure 4: Comparison of test accuracy across different submodel sizes for different submodel extraction methods on a global test dataset of CIFAR-10.
In Figure 4, we employ our algorithm as well as the baselines to extract submodels with different sizes and compare their performance on unparticipated clients. Generally, the performances of all methods increase as the size of extracted submodels grows. Our algorithm consistently outperforms the baseline methods. In contrast, alternative approaches, particularly static ones like HeteroFL and ScaleFL, suffer from more significant drops in performance. For the case of extracting a submodel that is much smaller than the minimum size involved in the training stage, our proposed method demonstrates remarkable superiority, outperforming existing works by at least 10%.
@@ 7 Conclusion
This work introduces an algorithm for model-heterogeneity FL, named FIARSE, that utilizes importance-aware operation to extract various sizes of submodels. In detail, we utilize TCB-GD that is able to optimize the clients’ local parameters to reflect their importance levels. Subsequently, we provide a theoretical analysis and highlight that the proposed work can converge to a neighborhood of a stationary point at the rate of $O\left(1/\sqrt{T}\right)$, where $T$ is the number of communication rounds. Extensive experiments are conducted on ResNet-18 and Roberta-base that demonstrate the significant superiority of our proposed method over the state-of-the-art works.
@@
This work addresses model heterogeneity in federated learning due to varying computational capacities among clients. The proposed method enhances the efficiency of on-device training and reduces computation and energy demands, which is particularly significant for resource-constrained devices like smartphones in real-world applications. Moreover, the method facilitates the practical deployment of federated learning systems in heterogeneous environments, making them more accessible and scalable.
Acknowledgements
The authors would like to thank the anonymous reviewers for their constructive comments. This work is supported in part by the US National Science Foundation under grants NSF IIS-1747614, NSF IIS-2226108, and NSF CNS-2154059. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.